OpenStudy (anonymous):

Find the indefinite integral sinø/sqrt(1-cosø)dø

OpenStudy (anonymous):

I used integral x^n dx = x^(n+1)(n+1)^-1+c and plugged the problem in to it and got 2sqrt(1-cosø)+c, but what happens to the sinø?

OpenStudy (jtvatsim):

is the question \[\int\limits_{}^{} \frac{\sin \theta}{ \sqrt{1- \cos \theta}} d \theta\] or is it \[\int\limits \frac{\sin \theta}{\sqrt{1 - \cos^2 \theta}} d \theta\]

OpenStudy (anonymous):

It is the first one

OpenStudy (jtvatsim):

Alright

OpenStudy (jtvatsim):

I don't think you can solve this by plugging it into the integral of x^n. Instead I think you need to use U-substitution.

OpenStudy (anonymous):

okay, I'll give that a try

OpenStudy (anonymous):